### v3

# Streamlined Supergravity

Abstract
The textbook N=1 supergravity has an F-term potential depending on a superpotential $W(z_{i})$ and a Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$, with the scalar potential $V(z^{i},\bar{z}^{\bar{i}})=e^{K}(|DW|^{2}-3|W|^{2})$. In this approach, it is not always easy to find the potential $V(z^{i},\bar{z}^{\bar{i}})$ with the required properties. We show that in supergravity with a nilpotent superfield and with any Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$ one can obtain any desired potential $V(z^{i},\bar{z}^{\bar{i}})$ by a proper choice of the Kähler metric of the nilpotent superfield. This construction is particularly suitable for cosmological and particle physics applications, which may require maximal freedom in the choice of kinetic terms and scalar potentials.
1 Introduction
Supergravity textbooks Wess:1992cp ; Freedman:2012zz and a more recent “Simple Supergravity” review DallAgata:2022nfr describe supergravities with linearly realized supersymmetry. The models with non-linearly realized supersymmetry have been actively studied over the last couple of decades, particularly in the context of cosmological applications of supergravity. These are best described as supergravity models that, in addition to unconstrained physical chiral multiplets $(Z^{i},\bar{Z}^{\bar{i}})$, contain one nilpotent chiral multiplet $(S,\bar{S})$ with $S^{2}=0$. This is a supergravity version of the anti-D3 brane uplifting of the potential in the KKLT model Kachru:2003aw required to produce de Sitter vacua in string theory. An early description of these models for inflation and dark energy can be found in Ferrara:2014kva , and in a more recent review Antoniadis:2024hvw .
The presence of a nilpotent superfield made it possible to construct pure de Sitter supergravity Bergshoeff:2015tra ; Hasegawa:2015bza , to stabilize all scalars but one Ferrara:2014kva ; Carrasco:2015rva , and to develop a large class of inflationary models currently favored by the data.
An important advantage of having the nilpotent multiplet in addition to standard chiral multiplets in the theory was emphasized in Kallosh:2015sea . The potential of these models at the vanishing scalar $s$ of the nilpotent multiplet has the form
$V=e^{K}(|F|^{2}+|DW|^{2}-3|W|^{2})\ ,$ (1)
where $|DW|^{2}-3|W|^{2}$ is the standard chiral superfield contributions to the potential. Meanwhile, the term $|F|^{2}$ is defined in eq. (2.6) in Kallosh:2015sea , so that
$V|_{s=0}=e^{K}(F_{s}K^{s\bar{s}}\bar{F}_{\bar{s}}+D_{i}WK^{i\bar{k}}\bar{D}_{\bar{k}}\bar{W}-3W\bar{W})\ .$ (2)
Here $D_{i}W$ is a Kähler covariant derivative of the superpotential over the scalar $z^{i}$ of the physical chiral multiplet $Z^{i}$, $F_{s}=D_{s}W$ where $D_{s}W$ is a Kähler covariant derivative of the superpotential over the scalar $s$ of the nilpotent chiral superfield multiplet $S$. Also, $K^{s\bar{s}}$ is the inverse Kähler metric of the scalar from a nilpotent superfield, and $K^{i\bar{k}}$ is the inverse Kähler metric of the physical scalars. Eq. (2) is valid under assumption made in Kallosh:2015sea that $K_{s}|_{s=0}=0$, in particular $K_{s\bar{k}}=0$.
Important steps in developing supergravity with one nilpotent multiplet in addition to standard chiral multiplets where made over the years in Kallosh:2015tea ; Kallosh:2015pho ; Ferrara:2016een ; McDonough:2016der ; Freedman:2017obq ; Kallosh:2017wnt ; Achucarro:2017ing ; Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha ; Kallosh:2025jsb . A related new construction called “Liberated Supergravity” was developed in Farakos:2018sgq based on linearly realized off-shell supersymmetry and Kähler - Weyl invariance.
Kähler manifolds with bisectional curvature along the $S=0$ plane in models with heavy stabilizer field $S$ were studied in Kallosh:2010xz , and with a nilpotent superfield $S^{2}=0$ in Kallosh:2017wnt :
$R^{\rm bisec}=-R_{z\,\bar{z}\,s\,\bar{s}}K^{z\,\bar{z}}K^{s\bar{s}}\,,\qquad R_{z\,\bar{z}\,s\,\bar{s}}=\partial_{z}\partial_{\bar{z}}K_{\bar{s}s}+K^{s\bar{s}}\partial_{z}K_{s\bar{s}}\partial_{\bar{z}}K_{s\bar{s}}\ .$ (3)
Kähler potentials of the form $K_{s\bar{s}}(z,\bar{z})S\bar{S}$ have a bisectional curvature iff $K_{s\bar{s}}$ depends on physical moduli $(z,\bar{z})$. It was shown in Kallosh:2010xz that terms involving the bisectional curvature play an important role in stabilizing moduli during inflation and in the properties of their inflationary fluctuations. Cosmological supergravity models with Kähler potentials mixing the stabilizer with a chiral sector $(z,\bar{z})$ were proposed in Kallosh:2010xz , and the ones mixing a nilpotent superfield with a chiral sector were proposed in McDonough:2016der ; Kallosh:2017wnt .
Now we take the main step in constructing streamlined supergravity: we express the nilpotent superfield Kähler metric via an arbitrary potential $V$ depending on the physical scalars, and an arbitrary Kähler potential $K$ of the physical scalars, using eq. (2). We bring this equation from Kallosh:2015sea to the form
$\boxed{K^{s\bar{s}}={e^{-K}V-D_{i}WK^{i\bar{k}}\bar{D}_{\bar{k}}\bar{W}-3W\bar{W}\over F_{s}\bar{F}_{\bar{s}}}\ .}$ (4)
We have thus multiplied eq. (2) by $e^{-K}\over F_{s}\bar{F}_{\bar{s}}$} and moved $K^{s\bar{s}}$ to the left and $e^{-K}V\over F_{s}\bar{F}_{\bar{s}}$} to the right. The resulting eq. (4) is now valid for any $V(z^{i},\bar{z}^{\bar{i}})$ and $K(z^{i},\bar{z}^{\bar{i}})$ depending on physical multiplets, same as the original equation (2).
Many examples of supergravities with nilpotent multiplets presented before in Kallosh:2015tea ; Kallosh:2015pho ; Ferrara:2016een ; McDonough:2016der ; Freedman:2017obq ; Kallosh:2017wnt ; Achucarro:2017ing ; Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha ; Kallosh:2025jsb , see also Kallosh:2022ggf ; Braglia:2022phb ; Kallosh:2024ymt ; Carrasco:2025rud , were restricted either to potentials at the inflationary trajectory, such as $V|_{z=\bar{z}}$, or to specific choices of the Kähler potentials of physical scalars. Now we have a general case of $V(z^{i},\bar{z}^{\bar{i}})$ and $K(z^{i},\bar{z}^{\bar{i}})$, explaining a posteriori why things were working so well in various examples.
We will present below a simple version of supergravity defined in Eqs. (2), (4) which achieves the goal of developing consistent supergravity with a nilpotent superfield, such that for any choice of the Kähler potential of the chiral superfields $K(z^{i},\bar{z}^{\bar{i}})$ one can obtain any desired potential $V(z^{i},\bar{z}^{\bar{i}})$ by a proper choice of the Kähler metric $K_{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})$ of the nilpotent superfield.
This is very different from the textbook N=1 supergravity with a potential depending on a superpotential $W(z_{i})$ and Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$ of the form $V_{\rm standard}(z^{i},\bar{z}^{\bar{i}})=e^{K}(|DW|^{2}-3|W|^{2})$, which often makes it difficult to obtain the desired potential $V(z^{i},\bar{z}^{\bar{i}})$ by a proper choice of $W(z_{i})$ and $K(z^{i},\bar{z}^{\bar{i}})$.
2 Construction of streamlined supergravity with a nilpotent superfield
Streamlining, by definition, means making something more efficient and simple by removing unnecessary parts. Note that the nilpotent sector acts behind the scenes; our only goal is to find a simple and efficient way to construct the most general scalar potentials and Kähler potentials in the physical sector. In our case, it involves two simplifications in the nilpotent sector of the theory.
  1. 1.
    Streamlining the dependence of the Kähler potential on the nilpotent superfield
    This step was already taken in Kallosh:2015sea ; Kallosh:2015tea , where the Kähler potential was taken in the form
    $K(z^{i},\bar{z}^{\bar{i}},s,\bar{s})=K(z^{i},\bar{z}^{\bar{i}})+K_{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})\,s\,\bar{s}\ .$ (5)
    A more general case would be the one where we add the $z,\bar{z}$-dependent terms that are linear in $s$ and $\bar{s}$. The result is known Schillo:2015ssx , supergravity is more complicated. But for our purposes, such complications are not warranted: we can reach our main goal without adding these extra terms in the nilpotent sector of the theory.
  2. 2.
    Streamlining the dependence of the superpotential $W$ on the nilpotent superfield, and making it $z$-independent
    We will use a simple superpotential
    $W=W_{0}+s\,F_{s}\ .$ (6)
The constant $W_{0}$ defines the gravitino mass, whereas the constant $F_{s}$ is a non-vanishing value of the $D_{s}W|_{s=0}=F_{s}$. We keep here a simple Kähler potential (5) and a simple superpotential (6), but our Kähler potential of physical scalars $K(z^{i},\bar{z}^{\bar{i}})$ is arbitrary. This choice of the simple $s$-sector is convenient and sufficient to construct supergravity with any desired potentials $V(z^{i},\bar{z}^{\bar{i}})$. Our choice of (6) simplifies the general expression for $F_{s}=D_{s}W$ used in Kallosh:2015sea and makes it possible to use a constant $F_{s}$ instead of a function $F_{s}(z^{i})$.
In case of global supersymmetry we have $n$ chiral superfields $Z^{i}=z^{i}+\sqrt{2}\,\theta\psi^{i}_{z}+\theta^{2}f_{z}^{i}$ and one nilpotent superfield $S$, where $S=s+\sqrt{2}\,\theta\psi^{s}+\theta^{2}f_{s}$ and $S^{2}=0$. The nilpotency condition is resolved as follows for $f_{s}\neq 0$ Komargodski:2009rz
$s={(\psi^{s})^{2}\over 2f_{s}}\ .$ (7)
This means that we have to impose a condition $s=0$ in the bosonic part of the action.
In the case of local supersymmetry, the resolution of the nilpotency condition is a generalization of the one in global supersymmetry. Superconformal theory with a nilpotent multiplet was studied in Kallosh:2015sea ; Kallosh:2015tea . Transition to supergravity followed the steps in Freedman:2012zz for usual chiral and vector multiplets, and special attention was given to the presence of the nilpotent chiral multiplet. As a result, the resolution of the nilpotency condition was established in Kallosh:2015sea ; Kallosh:2015tea and it takes the following form
$s={(\psi^{s})^{2}\over 2F}\,,\qquad{\rm where}\quad F=e^{K/2}K^{s\bar{s}}\,\big(-\overline{D_{s}W}+\dots\big)\ .$ (8)
Here $\dots$ denote various terms containing fermions. We will present more details in the Appendix to describe the complete supergravity action, including scalars, fermions, and vectors. In supergravity in Kallosh:2015sea and here in eq. (2) in absence of fermions
$\bar{F}_{\bar{s}}=\overline{D_{s}W}=-e^{-K/2}K_{s\bar{s}}\,F\ .$ (9)
With our choice of the streamlined nilpotent sector of the theory we need to identify the function $K_{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})$ or its inverse $K^{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})$ such that the standard supergravity expression for all chiral superfields, including the nilpotent one, in eq. (2), is reduced to a given arbitrary bosonic potential depending on $n$ complex scalars $V(z^{i},\bar{z}^{\bar{i}})$ when $s=0,\,F_{s}\neq 0$. The simplest way to do it is to replace the general $F_{s}(z^{i})=D_{s}W$ in eq. (4) by a constant $F_{s}$ for our choice of $W=W_{0}+s\,F_{s}$ and to take into account that $D_{i}W=K_{i}W_{0}$. This gives the following expression for the inverse Kähler metric:
$K^{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})={e^{-K(z^{i},\bar{z}^{\bar{i}})}V(z^{i},\bar{z}^{\bar{i}})+|W_{0}|^{2}(3-K_{i}K^{i\bar{k}}\bar{K}_{\bar{k}})\over F_{s}\bar{F}_{\bar{s}}}\ .$ (10)
The Kähler metric is identified as
$K_{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})={F_{s}\bar{F}_{\bar{s}}\over e^{-K(z^{i},\bar{z}^{\bar{i}})}V(z^{i},\bar{z}^{\bar{i}})+|W_{0}|^{2}(3-K_{i}K^{i\bar{k}}\bar{K}_{\bar{k}})}\ .$ (11)
Thus, the Kähler metric in streamlined supergravity depends only on $K(z^{i},\bar{z}^{\bar{i}})$ and its derivatives, and on $V(z^{i},\bar{z}^{\bar{i}})$.
To summarize: the streamlined supergravity depends on $n$ physical chiral superfields and one nilpotent superfield and has the following Kähler potential and superpotential
$K(z^{i},\bar{z}^{\bar{i}},s,\bar{s})=K(z^{i},\bar{z}^{\bar{i}})+{F_{s}\bar{F}_{\bar{s}}\over e^{-K(z^{i},\bar{z}^{\bar{i}})}V(z^{i},\bar{z}^{\bar{i}})+|W_{0}|^{2}(3-K^{i\bar{k}}K_{i}K_{\bar{k}})}\ s\,\bar{s}\,,\quad W=W_{0}+F_{s}\,s\ .$ (12)
We will sometimes call it “new” rather than “streamlined” for simplicity. The resulting bosonic action of scalars is
${{\cal L}(z^{i},\bar{z}^{\bar{i}})\over\sqrt{-g}}={R\over 2}+K_{i\bar{k}}(z^{i},\bar{z}^{\bar{i}})\,\partial z^{i}\partial\bar{z}^{\bar{k}}+K_{s\bar{s}}(z^{i},\bar{z}^{\bar{i}})\,\partial s\,\partial\bar{s}-V(z^{i},\bar{z}^{\bar{i}})\ ,$ (13)
where the Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$ and potential $V(z^{i},\bar{z}^{\bar{i}})$ are arbitrary.
The full supergravity action, including the fermions, was presented in the unitary gauge for local supersymmetry $\psi^{s}=0$ in Kallosh:2015sea , and a locally supersymmetric theory prior to gauge-fixing with a most general superpotential was given in Kallosh:2015tea . In Appendix A, we will give a short summary of the complete streamlined supergravity with scalars, fermions, and vectors following Kallosh:2015tea . In Appendix B, we will describe the implementation of our methods to $SL(2,\mathbb{Z})$ invariant supergravity.
3 Consistency issue of streamlined supergravity and its resolution
The issue of consistency was raised in a particular example of streamlined/new supergravity in Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha where it was observed that the expression for the Kähler metric $K_{s\bar{s}}$ of the kind we have in the general case in (11) for certain values of model parameters is not positive definite at some points in the moduli space. This was revealed for some choice of cosmological parameters in the examples studied in these papers. If $S$ were a normal chiral multiplet, one would have to require that $K_{s\bar{s}}>0$ to avoid a negative kinetic term for scalars and instability of the theory towards an infinite production of the corresponding particles.
An example given in Kallosh:2022vha is for a model with one chiral multiplet and one nilpotent multiplet:
$K(T,\bar{T},s,\bar{s})=-3\alpha\ln(T+\bar{T})+\frac{F_{s}\bar{F}_{\bar{s}}}{(T+\bar{T})^{3\alpha}V(T,\bar{T})+3|W_{0}|^{2}(1-\alpha)}\,s\,\bar{s}\ ,$ (14)
In this case, the bosonic action following from this supersymmetric construction is
${{\cal L}(T,\bar{T})\over\sqrt{-g}}={R\over 2}-{3\alpha\over 4}\,{\partial T\partial\bar{T}\over({\rm Re}\,T)^{2}}-V(T,\bar{T})\ .$ (15)
The potential at some point in the moduli space $T_{min}$ may have a de Sitter minimum where it is equal to $\Lambda\ll 1$ and we find that at this point
$K_{s\bar{s}}|_{T_{min}}=\frac{F_{s}\bar{F}_{\bar{s}}}{2^{3\alpha}\Lambda+3|W_{0}|^{2}(1-\alpha)}\,.$ (16)
The reason for the possible inconsistency is the fact that at the minimum with a very small cosmological constant $\Lambda$ only models with $\alpha<1$ have a positive definite Kähler metric $K_{s\bar{s}}$ since
$K_{s\bar{s}}(T,\bar{T})|_{T_{min}}<0\qquad\rm{if}\quad\alpha>1\ .$ (17)
This was viewed for a while in Yamada:2018nsk ; Linde:2018hmx ; Kallosh:2022vha as a restriction on the Kähler curvature of the moduli space, in this case, with $\mathcal{R}_{K}=-{2\over 3\alpha}$, the restriction was $|\mathcal{R}_{K}|<{2\over 3}$.
More recently, this issue was revisited in Kallosh:2025jsb , taking into account that $S$ is a nilpotent chiral multiplet. Consider the full supergravity action, which includes the Kähler metric $K_{s\bar{s}}$, which could lead to an inconsistency because it is not positive definite. These are kinetic terms in the action of the form
$K_{s\bar{s}}(\partial_{\mu}s\partial^{\mu}\bar{s}+\bar{\psi}^{s}\gamma_{\mu}D^{\mu}\psi^{s})\ .$ (18)
The nilpotent multiplet in supergravity satisfies the constraint that the scalar depends on the square of the spinor field
$s={(\psi^{s})^{2}\over 2F}\ .$ (19)
This makes the 1st term in the equation (18) quartic in spinor fields
$K_{s\bar{s}}\partial_{\mu}s\partial^{\mu}\bar{s}=K_{s\bar{s}}\partial_{\mu}\,{(\psi^{s})^{2}\over 2F}\partial^{\mu}\,{(\bar{\psi}^{\bar{s}})^{2}\over 2F}\ .$ (20)
These terms do not require $K_{s\bar{s}}$ to be positive. However, the second term in eq. (18) is quadratic in spinors, and its presence in the action indicates a possible inconsistency of the theory, as always in the case of the negative Kähler metric.
However, we can use local supersymmetry to gauge-fix the fermion from the nilpotent multiplet to vanish,
$\psi^{s}=0\ .$ (21)
This is a unitary gauge introduced in Bergshoeff:2015tra in supergravity interacting with one nilpotent multiplet and in Kallosh:2015sea in supergravity interacting with a nilpotent multiplet and other chiral multiplets. In this gauge, the second term in eq. (18) is absent, so the problem is solved, both terms in eq. (18) are absent from the gauge-fixed action.
Alternatively, one can take a different unitary gauge $v=0$ where the goldstino $v$ is interacting with the gravitino
$\bar{\psi}^{\mu}\gamma^{\mu}v=\bar{\psi}^{\mu}\gamma^{\mu}\Big[{1\over\sqrt{2}}e^{K(z^{i},\bar{z}^{\bar{i}})/2}(\psi_{z}^{i}D_{i}W+\psi^{s}D_{s}W)\Big]\ .$ (22)
The requirement that goldstino $v$ is vanishing means in our new supergravity models that
$v={1\over\sqrt{2}}e^{K(z^{i},\bar{z}^{\bar{i}})/2}(\psi^{i}_{z}K_{i}W_{0}+\psi^{s}F_{s})=0\quad\Rightarrow\quad\psi^{s}=-{\psi^{i}_{z}K_{i}(z^{i},\bar{z}^{\bar{i}})W_{0}\over F_{s}}\ .$ (23)
In the gauge where $v=0$, the fermion $\psi_{s}$ is replaced by a nonlinear function of moduli $(z^{i},\bar{z}^{\bar{i}})$ and a fermion $\psi^{i}_{z}$ since $K_{i}(z^{i},\bar{z}^{\bar{i}})$ does not have a moduli-independent term which would come from a Kähler potential linear in moduli $z^{i}$. But there is no reason to include the term $K(z,\bar{z})\sim z+\bar{z}$, it will not even contribute to the metric $K_{z\bar{z}}$.
Therefore, the kinetic terms of the boson $s$ and the fermion $\psi^{s}$ with a negative $K_{s\bar{s}}$ are not harmful. There is no vacuum instability due to these fermions since their kinetic term is either absent in $\psi^{s}=0$ gauge, or involves terms with negative $K_{s\bar{s}}$ which are not quadratic in physical fermions since they contain coupling to moduli $z^{i}$. Thus the streamlined supergravity defined by $K_{s\bar{s}}$ in eqs. (11) is consistent even in the models where $K_{s\bar{s}}<0$ at some points in moduli space.
4 Examples of streamlined supergravity models
In all examples, we use a simple $z^{i}$-independent superpotential $W=W_{0}+s\,F_{s}$ in eq. (6) and present Kähler potentials and the bosonic action. The Kähler metric of the nilpotent field is always an example of the general expression in (11).
4.1 Models with hyperbolic geometry and an arbitrary potential
4.1.1 E-models
For models with an exponential approach to the plateau, consider the case of E-models of $\alpha$-attractors Ferrara:2013rsa ; Kallosh:2013yoa . In streamlined supergravity, these models are described by complex half-plane variables $T=e^{-\sqrt{2\over 3\alpha}\varphi}+i\,a$ in hyperbolic geometry with the nilpotent superfield $S$ with $S^{2}=0$. The simplest models of this class have potentials with a plateau, which is approached exponentially fast at large $\varphi$. In particular, one may consider models (14), (15) with
$V_{E}(T,\bar{T})=\Lambda+V_{0}(1-T)^{n}(1-\bar{T})^{n}=\Lambda+V_{0}\Big(1-e^{-\sqrt{2\over 3\alpha}\varphi}\Big)^{2n}\ .$ (24)
Here $T=\bar{T}=e^{-\sqrt{2\over 3\alpha}\varphi}+i\,a$, and inflation occurs at large values of the canonically normalized field $\varphi$. To stabilize the axion field at $a=0$, one may simply add to this potential a term $m^{2}(T,\bar{T})(T-\bar{T})^{2}$, which vanishes for $T=\bar{T}$. (Note also that in this class of theories, one can construct consistent $\alpha$-attractor models with the same cosmological predictions without stabilizing the axion potential Achucarro:2017ing ; Linde:2018hmx ).
These models have a potential with a plateau, which is approached exponentially fast at large $\varphi$.
$V_{E}(T,\bar{T})=V_{0}\Big(1-2n\,e^{-\sqrt{2\over 3\alpha}\varphi}+\dots\Big)\ .$ (25)
The factor 2n in the last equation can be absorbed by a redefinition (a shift) of the field $\varphi$. This model can also be interpreted as pole inflation Galante:2014ifa ; Terada:2016nqg , with the kinetic terms having a pole of second order at $T=\bar{T}\to 0$ (15).
4.1.2 T-models
One can consider inflationary models with hyperbolic geometry also in disk variables $Z$ related to half-plane variables $T$ by a Cayley transform, as shown in Cecotti:2014ipa ; Kallosh:2015zsa
$T={1+Z\over 1-Z}\,,\qquad Z={T-1\over T+1}\ .$ (26)
One can show that kinetic terms of chiral superfields $T$ and $Z$ are exactly the same, up to a Cayley transform in (26)
${3\alpha\over 4}\,{\partial T\partial\bar{T}\over({\rm Re}\,T)^{2}}={3\alpha}\,{\partial Z\partial\bar{Z}\over(1-Z\bar{Z})^{2}}\ .$ (27)
Note that these kinetic terms have a pole of the order $p=2$.
The streamlined supergravity version of $\alpha$-attractors in disk variables, based on the general equation (4), corresponds to
$K(Z,\bar{Z},s,\bar{s})=$ $-3\alpha\ln(1-Z\bar{Z})+\frac{F_{s}\bar{F}_{\bar{s}}}{(1-Z\bar{Z})^{3\alpha}\,V(Z,\bar{Z})+3|W_{0}|^{2}(1-\alpha Z\bar{Z})}s\bar{s}\ .$ (28)
The bosonic action following from this supersymmetric construction is
${{\cal L}(Z,\bar{Z})\over\sqrt{-g}}={R\over 2}-{3\alpha}\,{\partial Z\partial\bar{Z}\over(1-Z\bar{Z})^{2}}-V(Z,\bar{Z})\ .$ (29)
Simplest examples include T-models of $\alpha$-attractors with $Z=\tanh{\varphi+i\vartheta\over\sqrt{6\alpha}}$ Kallosh:2013yoa . The potential with a heavy stabilized axion is
$V(Z,\bar{Z})=\Lambda+V_{0}(Z\bar{Z})^{n}+m^{2}(Z,\bar{Z})\,(Z-\bar{Z})^{2}\ .$ (30)
The inflaton potential for the canonically normalized inflaton field $\varphi$ is the T-model potential
$V_{\rm T}=\Lambda+V_{0}\tanh^{2n}{\varphi\over\sqrt{6\alpha}}\ .$ (31)
We assume that $\Lambda\ll V_{0}$, so one can ignore $\Lambda$ during inflation. The potential in disk variables with the stabilized axion and $Z=\bar{Z}=\tanh{\varphi\over\sqrt{6\alpha}}$ during inflation at large $\varphi$ is
$V(Z,\bar{Z})=V_{0}\tanh^{2n}{\varphi\over\sqrt{6\alpha}}=V_{0}\Big(1-2n\,e^{-\sqrt{2\over 3\alpha}\varphi}+\dots\Big)\ .$ (32)
We find the same exponential approach to the plateau in terms of a canonical field $\varphi$ as in the E-model (25). The factor 2n in the last equation can be absorbed by a redefinition (a shift) of the field $\varphi$. This is a case of the pole inflation in disk variables, where the kinetic terms have a pole of the order $p=2$ at the boundary $Z\bar{Z}\to 1$.
The supergravity version of polynomial $\alpha$-attractors Kallosh:2022feu is also described in (14), (15), with the kinetic term with the pole of order $p=2$. The potential in this case has particular features near the pole, where, e.g., for $k=2,4$
$V_{\rm pole}^{\rm ln}(T,\bar{T})\to V_{0}\Big(1-\Big({2\over\ln T\bar{T}}\Big)^{k}+\dots\Big)=V_{0}\Big(1-\Big({\sqrt{2\over 3\alpha}\varphi}\Big)^{-k}+\dots\Big)\ .$ (33)
In this model, the plateau at large $\varphi$ is approached polynomially Kallosh:2022feu ; Kallosh:2025ijd .
[…]
4.2.1 Pole inflation in half-plane variables
We now describe pole inflation with kinetic terms that have a pole of order $p>2$. To do it, we will take the following function $K(T,\bar{T},s,\bar{s})$ defined in the general equation (4):
$K(T,\bar{T},s,\bar{s})={\mu^{2}\over 2}({T}\overline{T})^{2-p\over 2}+{F_{s}\bar{F}_{\bar{s}}\over{V}\left({T},\overline{{T}}\right)e^{-{\mu^{2}\over 2}(T\overline{T})^{2-p\over 2}}+W_{0}^{2}\left(3-{\mu^{2}\over 2}(T\overline{T})^{2-p\over 2}\right)}\,s\bar{s}$ (34)
When the axion is stabilized, the bosonic action at $T=\bar{T}$ following from this supersymmetric construction is given by
${{\cal L}(T)\over\sqrt{-g}}={R\over 2}-{a_{p}\over 2}\,{(\partial T)^{2}\over T^{p}}-V(T)\ ,\qquad a_{p}={\mu^{2}\over 4}(p-2)^{2}\ .$ (35)
We may consider potentials near the pole of the kinetic term at $T\to 0$ of the form
$V|_{T=\bar{T}\to 0}\to V_{0}(1-T+\dots)\ .$ (36)
This model describes pole inflation Galante:2014ifa ; Terada:2016nqg with $p>2$, where in the canonical variables
$V|_{\varphi\to\infty}\to V_{0}(1-\Big({\mu\over\varphi}\Big)^{k}+\dots)\ ,\qquad k={2\over p-2}>0\ .$ (37)
4.2.2 Pole inflation in disk variables
The streamlined supergravity version of pole inflation with pole order $p>2$ in disk variables can be described by
$K(Z,\bar{Z},s,\bar{s})=$ ${\mu^{2}\over 2}(1-Z\bar{Z})^{2-p}+\frac{F_{s}\bar{F}_{\bar{s}}}{e^{{\mu^{2}\over 2}(1-Z\bar{Z})^{2-p}}\,V(Z,\bar{Z})+|W_{0}|^{2}\Big(3-{\mu^{2}\over 2}{(p-2)Z\bar{Z}(1-Z\bar{Z})^{p-2}\over 1+(p-2)Z\bar{Z}}\Big)}s\bar{s}\ .$ (38)
The bosonic action following from this supersymmetric construction is
${{\cal L}(Z,\bar{Z})\over\sqrt{-g}}={R\over 2}-{A_{p}\over 2}\,{\partial Z\partial\bar{Z}\over(1-Z\bar{Z})^{p}}-V(Z,\bar{Z})\ ,\qquad A_{p}={\mu^{2}}(p-2)\big(1+(p-2)Z\bar{Z}\big)\ .$ (39)
This is the pole-inflation model in which the kinetic terms have a pole of order $p>2$ at the boundary $Z\bar{Z}\to 1$. The polynomial approach to the plateau in terms of the canonical field $\varphi$ is similar to the one in the model (34). Observational results of these inflationary models depend only on $p$, on the value of the potential $V(Z,\bar{Z})$ at $Z\bar{Z}\to 1$, and on the value of $A_{p}$ at $Z\bar{Z}\to 1$.
$a_{p}=A_{p}\Big|_{Z\bar{Z}\to 1}=\mu^{2}(p-2)(p-1)\ .$ (40)
5 Discussion
In this paper we presented an efficient and relatively simple version of the streamlined $N=1$ supergravity with chiral multiplets and a nilpotent multiplet $S^{2}=0$ with the general type Kähler potential $K(z^{i},\bar{z}^{\bar{i}})$ leading to any desired scalar potential $V(z^{i},\bar{z}^{\bar{i}})$, once the proper choice of the Kähler metric of the nilpotent multiplet is made. The total Kähler potential, superpotential, and the bosonic action of this new supergravity are given in eqs. (12), (13). We provided some examples illustrating our methods in applications to inflationary potentials.
In Appendix A, we present a complete streamlined supergravity including scalars, vectors, fermions, and one nilpotent chiral multiplet with a general superpotential $W(z^{i},s)=g(z^{i})+s\,f(z_{i})$ as constructed in Kallosh:2015tea . The vector part of the theory and a holomorphic gauge kinetic function $f_{AB}(z^{i},s)=f_{AB0}(z^{i})+s\,f_{ABs}(z^{i})$ and momentum maps ${\cal P}_{A}$ correspond to a standard supergravity, as given in the book Freedman:2012zz . We also have a comment there about the work in Cribiori:2017laj concerning Fayet-Iliopoulos terms in supergravity without gauged R-symmetry.
The major difference with the book Freedman:2012zz before gauge-fixing local supersymmetry is in the fermionic sector of the theory, as one can see in Eq. (43). In the unitary gauge where the fermion of the nilpotent multiplet $\psi^{s}$ vanishes, all extra terms vanish. The resulting theory is given in Eq. (45) for physical and nilpotent scalars, vectors, and fermions, where one has to put the nilpotent scalar to zero. This last step leaves the additional part in the scalar potential, $e^{K}F_{s}K^{s\bar{s}}\bar{F}_{\bar{s}}$, relative to the standard supergravity without the nilpotent multiplet. This function is an arbitrary function of $(z^{i},\bar{z}^{\bar{i}})$.
In Appendix B, we apply this general method to $SL(2,\mathbb{Z})$ invariant supergravity theories.
We conclude that in the framework of streamlined supergravity, it is easy to find de Sitter solutions, to stabilize moduli, and to embed in supergravity many phenomenologically interesting bosonic models with all possible sets of $K(z^{i},\bar{z}^{\bar{i}})$ and $V(z^{i},\bar{z}^{\bar{i}})$.
Acknowledgement
We acknowledge many stimulating discussions over the years on supergravity with nilpotent multiplets with I. Antoniadis, J.J. Carrasco, S. Ferrara, D. Freedman, E. McDonough, D. Roest, M. Scalisi, A. Van Proeyen, T. Wrase, and Y. Yamada. This work was supported by the Leinweber Institute for Theoretical Physics at Stanford and by NSF Grant PHY-2310429.
Appendix A Complete streamlined supergravity with scalars, fermions, and vectors
The explicit supergravity action describing the interaction of supergravity with an arbitrary number of chiral and vector multiplets, as well as one nilpotent chiral multiplet, was presented in Kallosh:2015tea . It is defined by the Kähler potential of the form given in eq. (5), by the superpotential
$W(z^{i},s)=g(z^{i})+s\,f(z_{i})\ ,$ (41)
by a holomorphic gauge kinetic function $f_{AB}$
$f_{AB}(z^{i},s)=f_{AB0}(z^{i})+s\,f_{ABs}(z^{i})\ ,$ (42)
[…]
The complete supergravity action for the choice of the Kähler potential in (5), choice of $W$ in (41), choice of $f_{AB}$ in (42) was given in Kallosh:2015tea in eqs. (2.7) - (2.8) where $e^{-1}{\cal L}^{\rm book}$ are presented in detail in equations (18.6) - (18.19) in the book Freedman:2012zz . We reproduce the result in our notation here (where $s=z^{1}$)
$e^{-1}{\cal L}_{\rm final}=\Big[e^{-1}{\cal L}^{\rm book}\Big]_{s=\frac{(\psi^{s})^{2}}{2F^{s}}}-\frac{(\psi^{s})^{2}\,\,}{4K_{s\bar{s}}(F^{s}\bar{F}^{\bar{s}})^{2}}\left|K_{s\bar{s}}\frac{\Box(\psi^{s})^{2}}{2F^{s}}+B^{s}\right|^{2}\,\,\,$ (43)
with
$F^{s}\equiv K^{s\bar{s}}\left(-{\rm e}^{\frac{K}{2}}\,\overline{W}_{\bar{s}}+{1\over 2}(\partial_{i}K_{s\bar{s}})\bar{\psi}^{s}\psi^{i}+\frac{1}{4}\bar{f}_{AB\bar{s}}\bar{\lambda}^{A}P_{R}\lambda^{B}\right)\,,\quad B^{s}\equiv e^{-1}{\delta{\cal L}^{\rm book}\over\delta\bar{s}}|_{s=\bar{s}=0}\,.$ (44)
Note that the last term in equation (43) already has the maximal power of the (undifferentiated) spinor $\psi^{s}$, so that we can drop all terms in $B^{s}$ that contain $\psi^{s}$ or $\bar{\psi}^{\bar{s}}$. Explicit expressions for $B^{s}$ for some examples are given in Kallosh:2015tea .
This complete action with chiral and vector multiplets is complicated before gauge-fixing local supersymmetry. It is complicated before gauge-fixing, even in the case of pure de Sitter supergravity Bergshoeff:2015tra . But in the unitary gauge with $\psi^{s}=0$, both the action of the pure de Sitter supergravity in Bergshoeff:2015tra as well as the action with chiral and vector multiplets become very simple. In our case here, we find in the unitary gauge with $\psi^{s}=0$
$e^{-1}{\cal L}_{\rm final}=\Big[e^{-1}{\cal L}^{\rm book}\Big]_{s=0,\,\psi^{s}=0}$ (45)
In particular, we have a standard potential for all chiral multiplets as well as a leftover $V^{\rm book}_{s,\bar{s}}$ from the presence of the nilpotent multiplet, since $D_{s}W=F_{s}\neq 0$
$V|_{s=0}=V^{\rm book}_{\rm standard}+V^{\rm book}_{s,\bar{s}}=e^{K}(D_{i}WK^{i\bar{k}}\bar{D}_{\bar{k}}\bar{W}-3W\bar{W})+e^{K}(F_{s}K^{s\bar{s}}\bar{F}_{\bar{s}})\ .$ (46)
Finally, we would like to note here that particular choice of $K_{s\bar{s}}$ and $F_{s}$ in our scalar potential in (2) $e^{K}|F_{s}|^{2}K^{s\bar{s}}$ reproduces the potential (1.1) in Cribiori:2017laj . Namely, with $K_{s\bar{s}}=e^{K/3},|F_{s}|^{2}={\xi^{2}\over 2}$, we find that our potential is equal to ${\xi^{2}\over 2}e^{2K\over 3}$. This potential was interpreted in Cribiori:2017laj in the context of the gauge-fixed superconformal theory as a Fayet-Iliopoulos term in supergravity without gauged R-symmetry. It was, however, also observed there that in the context of constrained multiplets, this potential originates from the superpotential $W=W(z)-{\xi\over\sqrt{2}}s$. This is in agreement with our term F-term potential $e^{K}|F_{s}|^{2}K^{s\bar{s}}$ with $K_{s\bar{s}}=e^{K/3}$, $|F_{s}|^{2}={\xi^{2}\over 2}$.
Appendix B $SL(2,\mathbb{Z})$ invariant supergravity
We consider supergravity action with a physical chiral superfield where its scalar $\tau=iT$ is a modulus of the upper half-plane, and there is a nilpotent superfield with a scalar $s$ and fermion $\psi^{s}$. We study the action, including fermions, gravitino, and the fermion partner of $T$, which we call $\psi^{T}$, in the unitary gauge $\psi^{s}=0$. According to eq. (45), this means we have to start with standard supergravity with 2 chiral superfields with scalars $T,s$, and take into account that the potential (46) differs from the standard supergravity potential with only one superfield $T$.
Using the streamlined supergravity models with $W=W_{0}+s\,F_{s}$, one can construct a consistent supergravity embedding of the models with any potential $V(z^{i},\bar{z}^{\bar{i}})$. This method made it possible to construct a large set of supergravity models with $SL(2,\mathbb{Z})$- invariant potentials; see Sect. 9 of Kallosh:2024ymt . The Kähler invariant functional $G$ there is
$G(T,\bar{T};s,\bar{s})=-3\alpha\log(T+\bar{T})+\log|W_{0}+s\,F_{s}|^{2}+G_{s\bar{s}}(T,\bar{T})s\bar{s}\ ,$ (47)
where
$G_{s\bar{s}}=K_{s\bar{s}}={|F_{s}|^{2}\over(T+\bar{T})^{3\alpha}V(T,\bar{T})+3|W_{0}|^{2}(1-\alpha)}\ .$ (48)
The corresponding bosonic action is
${{\cal L}(T,\bar{T})\over\sqrt{-g}}={R\over 2}-{3\alpha\over 4}{\partial T\partial\bar{T}\over({\rm Re}T)^{2}}-V(T,\bar{T})\ .$ (49)
Here, the scalar kinetic term is $SL(2,\mathbb{Z})$ invariant. Therefore, if the potential $V(T,\bar{T})$ is $SL(2,\mathbb{Z})$ invariant, the total bosonic part of the supergravity action is $SL(2,\mathbb{Z})$ invariant. It was sufficient here to use the simplest superpotential of the streamlined supergravity $W=W_{0}+s\,F_{s}$.
However, if we also want to have the fermion part of the action with gravitino and $\psi^{T}$ to be $SL(2,\mathbb{Z})$ invariant, we should use a more general $T$-dependent superpotential of the kind given in eq. (41). In supergravity with one chiral multiplet without the nilpotent one, the $SL(2,\mathbb{Z})$ invariant theory, including gravitino and $\psi^{T}$ was presented in Ferrara:1989bc . The proof of $SL(2,\mathbb{Z})$ invariance there required that the generating functional
$G(T,\bar{T})=-3\alpha\log(T+\bar{T})+\log W(T)+\log\bar{W}(\bar{T})$ (50)
should be invariant under the $SL(2,\mathbb{Z})$ transformations. The superpotential $W(T)$ must transform as a modular form of weight $3\alpha$ to compensate the $SL(2,\mathbb{Z})$ transformation of $K(T,\bar{T})=-3\alpha\log(T+\bar{T})$. The gravitino, which is neutral with respect to the $SL(2,\mathbb{Z})$ transformations, couples to $G(T,\bar{T})$.
$e^{G(T,\bar{T})\over 2}\bar{\psi}_{\mu R}\sigma^{\mu\nu}\psi_{\nu R}+h.c.$ (51)
The fermion superpartner of $T$, the field $\psi^{T}$, transforms as a modular form of weight 2-2 and it couples to $G(T,\bar{T})$ and its derivatives. This is why the $SL(2,\mathbb{Z})$ invariance of $G(T,\bar{T})$ is necessary and sufficient for the invariance of the fermion part of the action in the case studied in Ferrara:1989bc .
Now, as different from Ferrara:1989bc , we start the action with two multiplets: the physical one and the nilpotent one. Gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})$ and its $T$ derivatives, where $G(T,\bar{T};s,\bar{s})=G(T,\bar{T})+\dots$ Here $\dots$ means terms linear in $s$ and $\bar{s}$ and in $s\bar{s}$. But according to eq. (45) in the unitary gauge $\psi^{s}=0$, we have to take the standard action for both multiplets at $\psi^{s}=0,s=0$. Thus, in the final action, gravitino and $\psi^{T}$ are coupled to $G(T,\bar{T};s,\bar{s})\Big|_{s=0}=G(T,\bar{T})$ and its derivatives in $T,\bar{T}$ as shown in eq. (7) in Ferrara:1989bc .
To construct this $SL(2,\mathbb{Z})$ invariant supergravity with one physical multiplet and one nilpotent multiplet, we proceed as follows. We choose a superpotential
$W(T,s)=W(T)+s\,F_{s}\ ,$ (52)
where $W(T)$ is a modular form of weight $3\alpha$. After that, our generating potential (50) becomes $SL(2,\mathbb{Z})$ invariant. In this superpotential, the constant $W_{0}$ is replaced by a function $W(T)$, but we still keep the constant $F_{s}$. Our generating functional is
$G(T,\bar{T};s,\bar{s})=-3\alpha\log(T+\bar{T})+\log(W(T)+sF_{s})+\log(\bar{W}(\bar{T})+\bar{s}\bar{F}_{s})+K_{s\bar{s}}(T,\bar{T})s\bar{s}$ (53)
As we explained above, in the unitary gauge the physical fermions couple only to $G(T,\bar{T};s,\bar{s}))\Big|_{s=0}$, which is an expression given in eq. (50). Therefore, the total supergravity action, bosonic and fermionic part, is $SL(2,\mathbb{Z})$ invariant under the condition that the potential $V(T,\bar{T})$ is $SL(2,\mathbb{Z})$ invariant and eq. (46) is satisfied. We therefore find that for any choice of the $SL(2,\mathbb{Z})$ invariant potential $V(T,\bar{T})$ the inverse metric of the nilpotent superfield is
$K^{s\bar{s}}(T,\bar{T})={(T+\bar{T})^{3\alpha}V(T,\bar{T})-D_{T}W(T)K^{T\bar{T}}\bar{D}_{\bar{T}}\bar{W}(\bar{T})+3W(T)\bar{W}(\bar{T})\over F_{s}\bar{F}_{\bar{s}}}\ .$ (54)
The total Kähler potential is
$K(T,\bar{T},s,\bar{s})=-3\alpha\log(T+\bar{T})+{F_{s}\bar{F}_{\bar{s}}\over(T+\bar{T})^{3\alpha}V(T,\bar{T})-D_{T}W(T)K^{T,\bar{T}}\bar{D}_{\bar{T}}\bar{W}(\bar{T})+3W(T)\bar{W}(\bar{T})}s\,\bar{s}$
and the total superpotential is given in Eq. (52). These expressions for $K(T,\bar{T},s,\bar{s})$ and $W(T,s)$ define the supergravity theory, which is $SL(2,\mathbb{Z})$ invariant.